# Enhancing Prompt Injection Attacks to LLMs via Poisoning Alignment

### v2

@@ 4.1 Experimental Setup
  • Accuracy:  Given an LLM $f$, we measure its standard accuracy on the well-known benchmarks MMLU [27], GPQA [28], and GSM8K [29].
As each $D_{t}$ or $D_{e}$ contains 100 samples, there are 10,000 pairs of samples when calculating ASVsoft or ASVhard for each target-injected task pair. To save computation cost, we randomly sample 100 pairs to compute ASVsoft or ASVhard following Liu et al. [7]. Additionally, when the target and injected tasks are identical, the injected task sample and target task sample have different ground-truth responses $r_{t}$ and $r_{e}$ in the 100 pairs. Unless otherwise stated, we report results from a single run, as we demonstrate that running multiple trials does not affect the results in Section 4.3. All experiments are run using 8 Nvidia Quadro-RTX-6000 GPUs. On average, aligning an LLM takes around 0.7 GPU-hours, and evaluating an LLM on the $7\times 7$ target-injected task pairs for one attack takes around 0.6 GPU-hours.
Table 2: Accuracy of LLMs on standard benchmarks after clean or poisoned alignment.
@@
Llama-2 Clean 0.462 0.161 0.227
Poisoned 0.464 0.172 0.232
Llama-3 Clean 0.634 0.297 0.766
Poisoned 0.636 0.306 0.753
Gemma Clean 0.508 0.209 0.303
Poisoned 0.510 0.202 0.301
Falcon Clean 0.264 0.121 0.033
Poisoned 0.259 0.148 0.037
Compared methods:  We extend two existing RLHF poisoning attacks [24, 22] to our scenario.
  • RLHFPoison [22]: : This method is a poisoning attack method designed to manipulate preference rankings in RLHF. It employs a three-step candidate selection process to stealthily alter human preference datasets by flipping the chosen samples’ preference labels while preserving alignment performance. In our extension, we create poisoned samples by flipping the preference labels of chosen samples and add them to the whole finetuning dataset.
  • Rando et al. [24]: : This method is a backdoor attack aiming to jailbreak aligned LLMs. It embeds a universal backdoor into the model by poisoning the RLHF training data. The backdoor trigger, a specific keyword, enables harmful responses universally, bypassing safety mechanisms without requiring prompt optimization. In our extension, following Rando et al. [24], for each poisoned sample we append “SUDO" at the end of target instruction as the backdoor trigger without concatenating with injection instruction, the construction of prefered response and rejected response keeps the same with PoisonedAlign.
4.2 Main Results
Table 3: ASV gap of Llama-3-8b-Instruct between poisoned and clean alignment under different poisoning methods. Rando et al.+ uses Combined Attack’s separator in Table 1 as the backdoor trigger.
Compared Methods ASV
@@ 4.2 Main Results
RLHFPoison [22] 0.01 -0.01
Rando et al. [24] -0.01 -0.02
@@
PoisonedAlign is effective:  For each LLM, we align it on poisoned or clean alignment data. Given a target-injected task pair, we compute the ASVsoft (or ASVhard) for the poisoned and clean LLMs. Then, we compute the ASVsoft (or ASVhard) gap between the poisoned and clean LLMs, and we average the gap over the $7\times 7$ target-injected task pairs. Table 0(b) shows such ASV gaps for each LLM and alignment dataset. Additional results on ASVhard and ASVsoft for each target-injected task pair before and after (poisoned or clean) alignment, are shown in Table 6 through Table 21 in Appendix. A higher ASV gap in Table 0(b) indicates an LLM becomes more vulnerable to prompt injection attacks due to poisoned alignment samples.
@@
Third, we find that Llama-3 and Gemma are more vulnerable to our PoisonedAlign compared to the other two LLMs. For example, on HH-RLHF, Llama-3 and Gemma respectively achieve ASVhard gaps of 0.27 and 0.32, while the other two LLMs have ASVhard gaps around 0.10. This increased vulnerability may be due to their stronger core instruction-following capabilities. Indeed, as shown in Table 2, these LLMs also perform better on standard benchmarks than the other two LLMs.
Table 4: ASV gap of an LLM between poisoned and clean alignment. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The alignment datasets are (a) HH-RLHF and (b) ORCA-DPO.
ASV Llama-2 Llama-3 Gemma Falcon
ASVhard 0.12 0.27 0.32 0.11
ASVsoft 0.07 0.11 0.18 0.10
ASV Llama-2 Llama-3 Gemma Falcon
ASVhard 0.08 0.33 0.26 0.06
ASVsoft 0.03 0.15 0.12 0.06
Figure 1: ASV gap of Llama-3 between poisoned and clean alignment for different $D$ and $D_{s}$. Each ASV gap is averaged over the $7\times 7$ target-injected task pairs.
Figure 2: Impact of (a) poisoning rate, (b) learning rate, and (c) epochs of DPO on PoisonedAlign.
@@
PoisonedAlign outperforms compared methods:  Table 3 presents the ASV gap between poisoned and clean alignment on Llama-3-8b-Instruct under default alignment settings. We have three key observations. First, our PoisonedAlign achieves a high ASV gap compared to all other methods. Second, RLHFPoison [22] and the default backdoor trigger from Rando et al. [24] have minimal impact on the ASV gap, with values close to 0. This suggests that these methods fail to make Llama-3-8b-Instruct more vulnerable to prompt injection. Third, using the Combined Attack’s separator as a backdoor trigger results in a higher ASVhard gap of 0.11 but a lower ASVsoft gap of 0.03. This highlights the importance of designing poisoned alignment samples that incorporate the injected instruction in the form of prompt injection rather than a conventional backdoor attack to effectively increase LLM vulnerability to prompt injection attacks.
@@
PoisonedAlign is stealthy:  The stealthiness goal ensures that an LLM retains its foundational capabilities after poisoned alignment, making PoisonedAlign difficult to detect based solely on an LLM’s performance in standard benchmarks. Table 2 shows the accuracy of LLMs on standard benchmarks after clean or poisoned alignment. In most cases, the accuracy gap between clean and poisoned alignment for an LLM is within 2%. This is because PoisonedAlign’s poisoned alignment samples remain high-quality, with preferred responses designed to follow the injected prompt’s instruction.
@@
Shadow dataset $D_{s}$ vs. alignment dataset $D$:  To generate poisoned alignment samples, our PoisonedAlign requires a shadow dataset $D_{s}$ with prompt-response pairs, which may or may not overlap with the alignment dataset $D$. Figure 1 shows the average ASV gap for Llama-3 between poisoned and clean alignment across the $7\times 7$ target-injected task pairs for different $D_{s}$ and $D$. We observe that PoisonedAlign remains effective whether or not $D_{s}$ overlaps with $D$. For example, when the alignment dataset is HH-RLHF, using $D_{s}$ from HH-RLHF (with overlap) and ORCA-DPO (without overlap) results in average ASVhard gaps of 0.27 and 0.31, and average ASVsoft gaps of 0.11 and 0.15, respectively. This is because PoisonedAlign crafts poisoned alignment samples without knowing the alignment dataset.
4.3 Ablation Study
Unless otherwise mentioned, in our ablation studies, we use Llama-3 as the LLM, HH-RLHF as the alignment dataset, hate detection as the target task, duplicate sentence detection as the injected task. More results on other target-injected task pairs can be found in Appendix.
@@ 4.3 Ablation Study
Table 5: ASV gap of an LLM between poisoned and clean alignment under supervised fine-tuning. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The datasets are (a) HH-RLHF and (b) ORCA-DPO.
ASV Llama-3 GPT-4o mini
ASVhard 0.09 0.20
ASVsoft 0.09 0.19
ASV Llama-3 GPT-4o mini
ASVhard 0.11 0.17
ASVsoft 0.10 0.16
PoisonedAlign is also effective for supervised fine-tuning:  Our main results are for preference alignment. For supervised fine-tuning, we report the ASV gaps between poisoned and clean LLMs in Table 2(e), averaged over the $7\times 7$ target-injected task pairs. We observe that our PoisonedAlign also achieves large ASVhard gap and ASVsoft gap on both Llama-3 and GPT-4o mini across alignment datasets. This is because PoisonedAlign crafts poisoned supervised fine-tuning data in a way that aligns an LLM to answer the injected prompts, making an LLM more vulnerable to prompt injection attacks after poisoned alignment.
Impact of poisoning rate:  The poisoning rate, defined as $|D^{\prime}|/|D|$ where $D^{\prime}$ is the set of poisoned samples, affects the performance of PoisonedAlign as shown in Figure LABEL:fig:poisoning_rate. ASVhard increases with the poisoning rate and converges beyond 0.1. Similar trends are observed across additional target-injected task pairs in Figure 3. This is because a higher number of poisoned alignment samples increases the likelihood of the LLM learning to complete injected prompts.
Figure 3: Impact of poisoning rate on PoisonedAlign for different target-injected task pairs.
Impact of alignment learning rate and epochs:  Figure LABEL:fig:learning_rate and Figure LABEL:fig:epochs illustrate the impact of the alignment learning rate and the number of DPO epochs on PoisonedAlign. Results for additional target-injected pairs can be found in Figure 4 and Figure 5. We observe that ASVhard is low when the learning rate is too small but stabilizes when the learning rate is within an appropriate range. For example, PoisonedAlign achieves below 0.40 ASVhard at a learning rate of $0.05\times 10^{-3}$, but achieves around 0.50 ASVhard when the learning rate is in the range of $[0.10,0.50]\times 10^{-3}$. This indicates that the LLM may be underfitting if the learning rate is too small. We also find that ASVhard initially increases and then converges as the number of alignment epochs increases. In particular, after a very small number of alignment epochs, an LLM becomes much more vulnerable to prompt injection attacks.
Figure 4: Impact of learning rate on PoisonedAlign for different target-injected task pairs.
Figure 5: Impact of epochs on PoisonedAlign for different target-injected task pairs.
Other prompt injection attacks:  By default, we use Combined Attack [7] for both crafting the poisoned alignment samples in PoisonedAlign and the evaluation in our experiments. Figure 6 shows the ASVhard for Llama-3 before and after clean/poisoned alignment, when the poisoned alignment samples are crafted using Combined Attack but the evaluation of the vulnerability uses other prompt injection attacks. Results for other target-injected task pairs are shown in Figure 7. We observe that the poisoned alignment data created by PoisonedAlign using Combined Attack also increases ASVhard for other prompt injection attacks after alignment, i.e., it makes the aligned LLM also more vulnerable to other attacks. This is because the poisoned alignment data generally aligns an LLM to complete injected prompts.
Figure 6: ASVhard for Llama-3 before and after clean/poisoned alignment on four additional prompt injection attacks: (a) Naive Attack, (b) Escape Characters, (c) Context Ignoring, and (d) Fake Completion.
@@
Figure 7: ASVhard for Llama-3 before and after clean/poisoned alignment on four additional prompt injection attacks: (a) Naive Attack, (b) Escape Characters, (c) Context Ignoring, and (d) Fake Completion. Injected task is hate detection and target task is spam detection.

### v3

@@ 4.1. Experimental Setup
  • Accuracy:  Given an LLM $f$, we measure its standard accuracy on the well-known benchmarks MMLU (Hendrycks et al., 2021), GPQA (Rein et al., 2024), and GSM8K (Cobbe et al., 2021).
As each $D_{t}$ or $D_{e}$ contains 100 samples, there are 10,000 pairs of samples when calculating ASVsoft or ASVhard for each target-injected task pair. To save computation cost, we randomly sample 100 pairs to compute ASVsoft or ASVhard following Liu et al. (2024a). Additionally, when the target and injected tasks are identical, the injected task sample and target task sample have different ground-truth responses $r_{t}$ and $r_{e}$ in the 100 pairs. Unless otherwise stated, we report results from a single run, as we demonstrate that running multiple trials does not affect the results in Section 4.3. All experiments are run using 8 Nvidia Quadro-RTX-6000 GPUs. On average, aligning an LLM takes around 0.7 GPU-hours, and evaluating an LLM on the $7\times 7$ target-injected task pairs for one attack takes around 0.6 GPU-hours.
Table 2. ASVs and corresponding gaps of an LLM between poisoned and clean alignment. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The alignment datasets are (a) $H$-RLHF and (b) ORCA-DPO.
(a) $H$-RLHF
ASV Alignment Llama-2-7b-chat Llama-3-8b-Instruct Gemma-7b-it Falcon-7b-instruct
ASVhard Clean 0.27 0.39 0.22 0.40
Poisoned 0.39 0.66 0.54 0.51
ASVhard Gap 0.12 0.27 0.32 0.11
ASVsoft Clean 0.42 0.55 0.38 0.43
Poisoned 0.49 0.66 0.56 0.53
ASVsoft Gap 0.07 0.11 0.18 0.10
(b) ORCA-DPO
ASV Alignment Llama-2-7b-chat Llama-3-8b-Instruct Gemma-7b-it Falcon-7b-instruct
ASVhard Clean 0.29 0.37 0.21 0.40
Poisoned 0.37 0.70 0.47 0.46
ASVhard Gap 0.08 0.33 0.26 0.06
ASVsoft Clean 0.42 0.56 0.38 0.41
Poisoned 0.45 0.71 0.50 0.47
ASVsoft Gap 0.03 0.15 0.12 0.06
Table 3. Accuracy of LLMs on standard benchmarks after clean or poisoned alignment.
@@
Llama-2-7b-chat
Clean
0.462 0.161 0.227
Poisoned
0.464 0.172 0.232
Llama-3-8b-Instruct
Clean
0.634 0.297 0.766
Poisoned
0.636 0.306 0.753
Gemma-7b-it
Clean
0.508 0.209 0.303
Poisoned
0.510 0.202 0.301
Falcon-7b-instruct
Clean
0.264 0.121 0.033
Poisoned
0.259 0.148 0.037
Compared methods:  We extend two existing RLHF poisoning attacks (Rando and Tramèr, 2023; Wang et al., 2024b) to our scenario.
  • RLHFPoison (Wang et al., 2024b): : This method is a poisoning attack method designed to manipulate preference rankings in RLHF. It employs a three-step candidate selection process to stealthily alter human preference datasets by flipping the chosen samples’ preference labels while preserving alignment performance. In our extension, we create poisoned samples by flipping the preference labels of chosen samples and add them to the whole finetuning dataset.
  • Rando and Tramèr (2023): : This method is a backdoor attack aiming to jailbreak aligned LLMs. It embeds a universal backdoor into the model by poisoning the RLHF training data. The backdoor trigger, a specific keyword, enables harmful responses universally, bypassing safety mechanisms without requiring prompt optimization. In our extension, following Rando and Tramèr (2023), for each poisoned sample we append “SUDO” at the end of target instruction as the backdoor trigger without concatenating with injection instruction, the construction of prefered response and rejected response keeps the same with PoisonedAlign.
4.2. Main Results
Table 4. ASV gap of Llama-3-8b-Instruct between poisoned and clean alignment under different poisoning methods. Rando et al.+ uses Combined Attack’s separator in Table 1 as the backdoor trigger.
Method ASV
@@ 4.2. Main Results
RLHFPoison (Wang et al., 2024b) 0.01 -0.01
 Rando and Tramèr (2023) -0.01 -0.02
@@
PoisonedAlign is effective:  For each LLM, we align it on poisoned or clean alignment data. Given a target-injected task pair, we compute the ASVsoft (or ASVhard) for the poisoned and clean LLMs. Then, we compute the ASVsoft (or ASVhard) gap between the poisoned and clean LLMs, and we average the gap over the $7\times 7$ target-injected task pairs. Table 2 shows such ASV gaps for each LLM and alignment dataset. A higher ASV gap in Table 2 indicates an LLM becomes more vulnerable to prompt injection attacks due to poisoned alignment samples. Figures 8, 9, 10 and 11 in Appendix show the detailed results for ASVhard and ASVsoft of different pairs of target and injected tasks on default LLM Llama-3-8b-Instruct finetuned on HH-RLHF (Bai et al., 2022) and ORCA-DPO (Intel, 2024).
@@
Third, we find that Llama-3-8b-Instruct and Gemma-7b-it are more vulnerable to our PoisonedAlign compared to the other two LLMs. For example, on HH-RLHF, Llama-3-8b-Instruct and Gemma-7b-it respectively achieve ASVhard gaps of 0.27 and 0.32, while the other two LLMs have ASVhard gaps around 0.10. This increased vulnerability may be due to their stronger core instruction-following capabilities. Indeed, as shown in Table 3, these LLMs also perform better on standard benchmarks than the other two LLMs.
@@
PoisonedAlign outperforms compared methods:  Table 4 presents the ASV gap between poisoned and clean alignment on Llama-3-8b-Instruct under default alignment settings. We have three key observations. First, our PoisonedAlign achieves a high ASV gap compared to all other methods. Second, RLHFPoison (Wang et al., 2024b) and the default backdoor trigger from  Rando and Tramèr (2023) have minimal impact on the ASV gap, with values close to 0. This suggests that these methods fail to make Llama-3-8b-Instruct more vulnerable to prompt injection. Third, using the Combined Attack’s separator as a backdoor trigger results in a higher ASVhard gap of 0.11 but a lower ASVsoft gap of 0.03. This highlights the importance of designing poisoned alignment samples that incorporate the injected instruction in the form of prompt injection rather than a conventional backdoor attack to effectively increase LLM vulnerability to prompt injection attacks.
@@
PoisonedAlign is stealthy:  The stealthiness goal ensures that an LLM retains its foundational capabilities after poisoned alignment, making PoisonedAlign difficult to detect based solely on an LLM’s performance in standard benchmarks. Table 3 shows the accuracy of LLMs on standard benchmarks after clean or poisoned alignment. In most cases, the accuracy gap between clean and poisoned alignment for an LLM is within 2% and even on GPQA benchmark we observe better performance for poisoned models consistently. This is because PoisonedAlign’s poisoned alignment samples remain high-quality, with preferred responses designed to follow the injected prompt’s instruction.
@@
Shadow dataset $D_{s}$ vs. alignment dataset $D$:  To generate poisoned alignment samples, our PoisonedAlign requires a shadow dataset $D_{s}$ with prompt-response pairs, which may or may not overlap with the alignment dataset $D$. Figure 2 shows the average ASV gap for Llama-3-8b-Instruct between poisoned and clean alignment across the $7\times 7$ target-injected task pairs for different $D_{s}$ and $D$. We observe that PoisonedAlign remains effective whether or not $D_{s}$ overlaps with $D$. For example, when the alignment dataset is HH-RLHF, using $D_{s}$ from HH-RLHF (with overlap) and ORCA-DPO (without overlap) results in average ASVhard gaps of 0.27 and 0.31, and average ASVsoft gaps of 0.11 and 0.15, respectively. This is because PoisonedAlign crafts poisoned alignment samples without knowing the alignment dataset.
(a) ASVhard
(b) ASVsoft
@@
Figure 2. ASV gap of Llama-3-8b-Instruct between poisoned and clean alignment for different $D$ and $D_{s}$. Each ASV gap is averaged over the $7\times 7$ target-injected task pairs.
[…]
Two heatmaps showing ASV gaps for Llama 3 under poisoned vs. clean alignment on hard and soft cases.
(a) Poisoning rate
(b) Learning rate
(c) Epochs
Figure 3. Impact of (a) poisoning rate, (b) learning rate, and (c) epochs of DPO on PoisonedAlign.
[…]
Three subfigures showing how poisoning rate, learning rate, and number of epochs affect DPO performance.
@@
Figure 4. ASVhard for Llama-3-8b-Instruct before and after clean/poisoned alignment on four additional prompt injection attacks: (a) Naive Attack, (b) Escape Characters, (c) Context Ignoring, and (d) Fake Completion.
[…]
Four subfigures showing ASVhard results for Llama 3 under different prompt injection attacks.
4.3. Ablation Study
Unless otherwise mentioned, in our ablation studies, we use Llama-3-8b-Instruct as the LLM, HH-RLHF as the alignment dataset, hate detection as the target task, duplicate sentence detection as the injected task. More results on other target-injected task pairs can be found in Appendix.
Table 5. ASVs and corresponding gaps of an LLM between poisoned and clean alignment under supervised fine-tuning. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The datasets are (a) $H$-RLHF and (b) ORCA-DPO.
(a) $H$-RLHF
ASV Alignment Llama-3-8b-Instruct GPT-4o mini
ASVhard Clean 0.45 0.40
Poisoned 0.54 0.60
ASVhard Gap 0.09 0.20
ASVsoft Clean 0.44 0.40
Poisoned 0.53 0.59
ASVsoft Gap 0.09 0.19
(b) ORCA-DPO
ASV Alignment Llama-3-8b-Instruct GPT-4o mini
ASVhard Clean 0.54 0.52
Poisoned 0.65 0.69
ASVhard Gap 0.11 0.17
ASVsoft Clean 0.56 0.53
Poisoned 0.66 0.69
ASVsoft Gap 0.10 0.16
PoisonedAlign is also effective for supervised fine-tuning:  Our main results are for preference alignment. For supervised fine-tuning, we report the ASV gaps between poisoned and clean LLMs in Table 5, averaged over the $7\times 7$ target-injected task pairs. We observe that our PoisonedAlign also achieves large ASVhard gap and ASVsoft gap on both Llama-3-8b-Instruct and GPT-4o mini across alignment datasets. This is because PoisonedAlign crafts poisoned supervised fine-tuning data in a way that aligns an LLM to answer the injected prompts, making an LLM more vulnerable to prompt injection attacks after poisoned alignment.
Impact of poisoning rate:  The poisoning rate, defined as $|D^{\prime}|/|D|$ where $D^{\prime}$ is the set of poisoned samples, affects the performance of PoisonedAlign as shown in Figure 3(a). ASVhard increases with the poisoning rate and converges beyond 0.1. Similar trends are observed across additional target-injected task pairs in Figure 7 in Appendix. This is because a higher number of poisoned alignment samples increases the likelihood of the LLM learning to complete injected prompts.
Figure 5. Impact of Llama-3-8b-Instruct’s temperature on PoisonedAlign.
[…]
Figure 6. Impact of the number of repeated trials on PoisonedAlign. Across all trials, the standard deviation consistently remains close to 0.
Impact of alignment learning rate and epochs:  Figure 3(b) and Figure 3(c) illustrate the impact of the alignment learning rate and the number of DPO epochs on PoisonedAlign. Results for additional target-injected pairs can be found in Figure 8 and Figure 9 in Appendix. We observe that ASVhard is low when the learning rate is too small but stabilizes when the learning rate is within an appropriate range. For example, PoisonedAlign achieves below 0.40 ASVhard at a learning rate of $0.05\times 10^{-3}$, but achieves around 0.50 ASVhard when the learning rate is in the range of $0.10,0.50]\times 10^{-3}$. This indicates that the LLM may be underfitting if the learning rate is too small. We also find that ASVhard initially increases and then converges as the number of alignment epochs increases. In particular, after a very small number of alignment epochs, an LLM becomes much more vulnerable to prompt injection attacks.
Other prompt injection attacks:  By default, we use Combined Attack (Liu et al., 2024a) for both crafting the poisoned alignment samples in PoisonedAlign and the evaluation in our experiments. Figure 4 shows the ASVhard for Llama-3-8b-Instruct before and after clean/poisoned alignment, when the poisoned alignment samples are crafted using Combined Attack but the evaluation of the vulnerability uses other prompt injection attacks. Results for other target-injected task pairs are shown in Figure 10 in Appendix. We observe that the poisoned alignment data created by PoisonedAlign using Combined Attack also increases ASVhard for other prompt injection attacks after alignment, i.e., it makes the aligned LLM also more vulnerable to other attacks. This is because the poisoned alignment data generally aligns an LLM to complete injected prompts.